Write these in rectangular form in x, y:

x(t)=t^2-1, y(t)=3t^3+2t
Answer #1

we have

x=t^2-1

x+1=t^2

t=(x+1)^\frac{1}{2}

and

y=3t^2+2t

y=3t^3+2t

put t = (x+1)1/2,

y=3(x+1)^\frac{3}{2}+2(x+1)^\frac{1}{2}

y=(x+1)^\frac{1}{2}(3x+3+2)

y=(x+1)^\frac{1}{2}(3x+5)